How do I subtrack 2 5/8 feet from 8 feet
Found 2 solutions by Shin123, josgarithmetic:
Answer by Shin123(626)

Convert 8 into a fraction, 64/8. Convert 2 5/8 into 21/8. 64/8-21/8=43/8=5 3/8.

Answer by josgarithmetic(39617)